4x^(2)-7x+3= can u show me the work for this problem and also the answer
Answer by checkley79(3341)

4x^(2)-7x+3=0
(4X-3)(X-1)=0
4X-3=0
4X=3
X=3/4 ANS.
X-1=0
X=1 ANS.
PROOF:
4*3/4^2-7*3/4+3=0
4*9/16-21/4+3=0
36/16-21/4+3=0
2.25-5.25+3=0
3=3
4*1^2-7*1+3=0
4-7+3=0
7-7=0
0=0